A famous golfer has finished
in the first 10 places in 22 out
of 39 tournaments so far this
year. Her sponsor says that she
must finish in the top 10 places
at least 60% of the time to
qualify for a bonus. If there are
16 tournaments yet to play, in
how many of these does she
need to finish in the top 10 in
order to earn her bonus?

Answer by Theo(13342)

A famous golfer has finished in the first 10 places in 22 out of 39 tournaments so far this year. Her sponsor says that she must finish in the top 10 places at least 60% of the time to qualify for a bonus. If there are 16 tournaments yet to play, in how many of these does she need to finish in the top 10 in order to earn her bonus?

she has played in 39 so far and she still have 16 more tournaments to play.
39 + 16 = 55 total tournaments.
60% of 55 = .6 * 55 = 33.

she needs to place in the top 10 in 33 of the 55 tournaments.
she has already placed in the top 10 for 22 of them.
she needs to place in the top 10 for 11 more.
22 + 11 = 33.